OpenStudy (anonymous):

2/4 4/5 6/6 8/7 ... if this pattern continues what number will belong here 100/?

geerky42 (geerky42):

\(\dfrac{2n}{n+3}\)

geerky42 (geerky42):

So 2n is equal to 100. find the value of n then plug that in denominator. hope that helps.

OpenStudy (anonymous):

i think i messed up and put 20 for my answer

geerky42 (geerky42):

ok, so do you understand what I have explained so far?

OpenStudy (anonymous):

yes , thanks

OpenStudy (anonymous):

2*100=200 2+3=5 , 200/5 = 40